OpenStudy (anonymous):

Solve -37 + n = -56 for n.

OpenStudy (confluxepic):

Add 37 on both sides.

TheSmartOne (thesmartone):

-37+n=-56 n-37=-56 +37 +37 n=-56+37 n=?

OpenStudy (confluxepic):

n=-56+37

TheSmartOne (thesmartone):

Or n=37-56

OpenStudy (confluxepic):

n=?`

OpenStudy (confluxepic):

@Ayo_CH33TO

OpenStudy (anonymous):

n= -19?

OpenStudy (confluxepic):

@TheSmartOne

OpenStudy (confluxepic):

Correct.

OpenStudy (confluxepic):

\[\huge n=-19\]

TheSmartOne (thesmartone):

YEs n=-19